OpenStudy (anonymous):

integral of 1/e^x dx

OpenStudy (hoblos):

1/e^x = e^(-x)

OpenStudy (hoblos):

can you do it now

OpenStudy (anonymous):

\[\frac{1}{e^{x}}=e^{-x}\]and then substitute\[u=-x\]

OpenStudy (ash2326):

@ncorrea92 did you understand the problem?

OpenStudy (amistre64):

why would we substitute u = -x?

OpenStudy (anonymous):

if you already know the antiderivative of e^(ax) then you dont need any substitution

OpenStudy (anonymous):

its just e^-x then ?

OpenStudy (hoblos):

well integral e^(ax) = (1/a)e^(ax) so here integral e^(-x) = -e^(-x)

OpenStudy (anonymous):

oh alright thanks

OpenStudy (hoblos):

yw =)